OpenStudy (hba):

If x – 1, x + 2 and x + 1 are all factors of a polynomial P(x) of degree 3, which of the following must also be a factor of P(x)? I) x^2+ 3x + 2 II) x^2– 1 III) x^2+ x – 2 A) III only B) II only C) I only D) All three are factors

OpenStudy (hba):

@dpaInc

OpenStudy (anonymous):

HINT: since 2, 3, 5 are factors of 30, then: 2*3 = 6 2*5 = 10 3*5 = 15 are also factors of 30.....

OpenStudy (anonymous):

in other words, since (x-1), (x+2), (x+1) are factors of P(x), then: (x-1)(x+2) = ??? (x-1)(x+1) = ??? (x+1)(x+2) = ??? are also factors of P(x)....